# Nonconvergence of the Feynman–Dyson diagrammatic perturbation expansion of propagators

### v7

@@ III.1 Formalisms
In most applications, both the Green’s function and self-energy are expanded in perturbation series. In this article, the first- and higher-order perturbation corrections are denoted by symbols prefixed with $\delta$ with its order given as the parenthesized superscript.
@@
where $\bm{G}(\omega;\lambda)$ and $\bm{\Sigma}(\omega;\lambda)$ are the exact (i.e., FCI) values of the respective quantities for a perturbation-scaled Hamiltonian $\hat{H}=\hat{H}^{(0)}+\lambda\hat{V}^{(1)}$. The zeroth-order Hamiltonian $\hat{H}^{(0)}$ corresponds to the $\bm{G}^{(0)}$ of Eq. (5). Here, we adopt the HF theory as the zeroth order, which implies $\delta{\Sigma}_{pq}^{(1)}={0}$ Hirata et al. (2017). In this article, MBGF($n$) refers to the $n$th-order Feynman–Dyson perturbation theory, whereas the exact MBGF method means FCI described in Sec. II. MBGF($n$) is diagrammatically linked Hirata et al. (2017) and size-consistent Hirata (2011), but the exact MBGF method is a non-size-consistent configuration interaction, and is thus not useful for larger molecules or solids.
These $\lambda$-derivatives can be taken either numerically or analytically. From the former, we obtain benchmark data of the perturbation corrections at several low orders Hirata et al. (2017). From the latter, we derive recursions of $\delta\bm{G}^{(n)}$ and $\delta\bm{\Sigma}^{(n)}$ in the style of Rayleigh–Schrödinger perturbation theory, which can then be implemented into a general-order algorithm Hirata et al. (2017). This algorithm was used in the present study. The recursions also justify the diagrammatic rules through the linked-diagram and irreducible-diagram theorems in a time-independent picture Hirata et al. (2017).
While the time-independent picture is more mathematically transparent and systematically extensible to arbitrarily high orders Hirata et al. (2017), the time-dependent one may be more intuitive and expedient Dyson (1993); Mattuck (1992). In the latter, for example, the second-order self-energy is stipulated diagrammatically as in Fig. 5. It graphically describes the process in which (i) a mean-field particle (hole) scatters another particle out of its mean-field state, thereby creating a hole, at one time; (ii) all three particles and holes propagate in their respective mean-field potentials, i.e., driven by the mean-field propagator $\bm{G}^{(0)}$; and (iii) the particle-hole pair recombines at another time. The numerical value of $\delta\bm{\Sigma}^{(2)}$, which is related to the probability of this overall process, is the product of the probabilities of the constituent scattering and propagation events summed over all possible times and positions of their occurrences. Consulting with Table 4.3 of Mattuck Mattuck (1992), we can then evaluate it as
@@
where an occupied spinorbital index and the imaginary unit both denoted by “$i$” may be easily distinguished, and $\langle pq||rs\rangle$ is an antisymmetrized two-electron integral Szabo and Ostlund (1982); Shavitt and Bartlett (2009). Second-order many-body Green’s-function method [MBGF(2)] solves the inverse Dyson equation [Eq. (9)] with this $\bm{\Sigma}^{(2)}$. Since the roots of this equation occur at the intersection of $\bm{\epsilon}+\bm{\Sigma}^{(2)}(\omega)$ and $\omega$, they cannot be divergent even though the method is perturbative.
@@
in analogy to Eq. (7), which is also consistent with the Dyson equations [Eqs. (4) and (4)]. Therefore, diagrammatically, as shown in Fig. 7, $\bm{G}^{\text{Dyson}(2)}$ is a bold-line Green’s function (just like the one appearing in Fig. 1), which includes an infinite-order correction through repeated actions of $\bm{\Sigma}^{(2)}$. In this sense, MBGF($n$) is an infinite-order theory for IPs and EAs even for a finite perturbation order $n$.
@@
corresponding to the diagrammatic equation in Fig. 8. It may be noticed that the functional form of $\delta\bm{\Sigma}^{(3)}$ with respect to $\omega$ is completely different from that of $\delta\bm{\Sigma}^{(2)}$ [Eq. (25)]; poles in $\delta\bm{\Sigma}^{(3)}$ are second order, while $\delta\bm{\Sigma}^{(2)}$ and the exact Green’s function have only first-order poles. This difference has a grave consequence on the roots of the inverse Dyson equation, which we now discuss.
@@ III.2 Numerical results
Figure 9: The diagonal ${\epsilon_{3}}+{\Sigma_{33}^{(2)}}(\omega)$ and ${\epsilon_{3}}+\Sigma_{33}^{(3)}(\omega)$ as functions of $\omega$ for the BH molecule.
@@
Figure 10: The diagonal $G_{33}^{\text{Dyson}(2)}(\omega)$ and $G_{33}^{\text{Dyson}(3)}(\omega)$ as functions of $\omega$ for the BH molecule.
@@
In Fig. 9 are plotted diagonal elements of the second- and third-order self-energies, $\epsilon_{3}+\Sigma_{33}^{(2)}$ and $\epsilon_{3}+\Sigma_{33}^{(3)}$, for the third orbital (HOMO) of the BH molecule Hirata et al. (2017). Intersections of these functions with the diagonal $\omega$ line (also drawn) are the roots of the corresponding inverse Dyson equations (in the diagonal approximation, in this case).
@@
The second-order self-energy has qualitatively the same functional form as the exact self-energy (Fig. 3) in that they are both separated by singularities into consecutive $\omega$ brackets, within each of which they are monotonically decreasing. The singularities of $\bm{\Sigma}^{(2)}$, i.e., the boundaries of the brackets, occur at the two-particle-one-hole (2p1h) ($\epsilon_{a}+\epsilon_{b}-\epsilon_{i}$) and two-hole-one-particle (2h1p) ($\epsilon_{i}+\epsilon_{j}-\epsilon_{a}$) HF orbital energy differences according to Eq. (25). In each bracket, a diagonal element of the second-order self-energy intersects the $\omega$ line exactly once, just as the exact self-energy does. The second-order inverse Dyson equation has a principal root for the HOMO at around $-0.25\,E_{\text{h}}$ in the central bracket enclosing $\omega=0$, where the self-energy is relatively flat and whose residue is close to unity [see Eq. (14)]. It also has one satellite root in each of the other brackets, where the self-energy is a near-vertical dive and whose residue is close to zero. That the self-energy is nearly vertical at these satellite roots implies that these roots practically coincide with the 2p1h or 2h1p HF orbital energy differences, accounting for little to no electron-correlation effects. This problem becomes even acuter at higher orders. Nevertheless, MBGF(2) is overall well behaved.
@@
The third-order self-energy, in contrast, has a qualitatively wrong functional form. It is still separated by the same 2p1h and 2h1p singularities into the identical set of brackets [see Eq. (27)], but within each bracket, $\bm{\Sigma}^{(3)}$ is either concave or convex except in the central bracket, where it is monotonically decreasing. Consequently, in the domain of $\omega$ of Fig. 9, the third-order inverse Dyson equation has no real roots except the principal root for the HOMO at around $-0.25\,E_{\text{h}}$, where the third-order self-energy closely traces the second-order self-energy. MBGF(3) is, therefore, qualitatively wrong, and also differs dramatically from MBGF(2) (except near $\omega=0$), even though they are two adjacent members of a hierarchical approximation series. The different functional forms of the self-energy between the second and third orders can be easily rationalized by their algebraic definitions, Eqs. (25) and (27). The second-order self-energy (as well as the exact self-energy) has only the first-order poles, whereas the third-order self-energy features up to the second-order poles.
@@
Figure 11: All six diagonal elements of $\bm{\epsilon}+\bm{\Sigma}^{(3)}(\omega)$ as functions of $\omega$ for the BH molecule. The fourth and fifth elements are degenerate, appearing as one (purple) curve. The 2p1h and 2h1p HF orbital energy differences are superposed.
Where have the missing satellite roots of MBGF(3) gone? We believe that they are in the complex space, although no attempt was made to determine them numerically.
@@
From Fig. 11, it can furthermore be observed that real satellite roots of MBGF(3) tend to appear as nearly degenerate pairs or doublets (which is distinct from the degeneracy of roots mentioned in the foregoing paragraphs). One of a doublet corresponds to an intersection of a sharply falling part of the self-energy with the $\omega$ line and thus has a positive infinitesimal residue, while the other intersection occurs at an almost vertically rising part of the self-energy, whose residue is negative (which is nonphysical). These intersections almost coincide with bare 2p1h or 2h1p HF orbital energy differences, thus accounting for little to no correlation effects.
[…]
The $\bm{\Sigma}^{(9)}$ has a concave or convex shape in each bracket separated by its singularities and, as a result, does not intersect the $\omega$ line as often as $\bm{\Sigma}^{(8)}$ does. In Fig. 12, the $\bm{\Sigma}^{(9)}$ intersects the $\omega$ line only once. In Fig. 12, the second through sixth diagonal elements of $\bm{\Sigma}^{(9)}$ cross the $\omega$ line numerous times in $-18\,E_{\text{h}}\leq\omega\leq-15\,E_{\text{h}}$, forming manifolds of roots many of whose residues are nonphysical. The first diagonal element of $\bm{\Sigma}^{(9)}$ has several crossings with the $\omega$ line in the same $\omega$ window and also in $-3\,E_{\text{h}}\leq\omega\leq-1\,E_{\text{h}}$. While these intersections are doublets, their occurrences are likely fewer than the intersections of $\bm{\Sigma}^{(8)}$, and many satellite roots of MBGF(9) are therefore believed to be complex.
@@
Our numerical tests for other molecules (not shown) have indicated that an odd-order self-energy can generally be concave, convex, or monotonically increasing within a bracket, while it is monotonically decreasing in the central bracket enclosing $\omega=0$. Therefore, an odd-order inverse Dyson equation can have (many) real roots (sometimes more than one in a bracket), but the corresponding residues may fall outside the valid range of zero to one.
The eighth-order self-energy is less problematic in this regard, but it comes with new issues. Its brackets are separated by the 2p1h, 2h1p, 3p2h, 3h2p, 4p3h, 4h3p, 5p4h, and 5h4p HF orbital energy differences. The $\bm{\Sigma}^{(8)}$ has extremely sharp singularities at these frequencies, where it intersects the $\omega$ line. Therefore, although MBGF(8) has only real roots, most of the satellite roots are no different from these bare HF orbital energy differences (superposed in Fig. 12), accounting for virtually no correlation effects.
Figure 14: The diagonal ${G_{33}}(\omega)$, $G_{33}^{\text{Dyson}(8)}(\omega)$, and $G_{33}^{\text{Dyson}(9)}(\omega)$ as functions of $\omega$ for the BH molecule. The exact IPs and EAs obtained by FCI are superposed.
Figure 14, which plots the eighth- and ninth-order bold-line Green’s functions, reiterates the above observations. The odd-order Green’s function $\bm{G}^{\text{Dyson}(9)}$ shows much fewer poles, often displaying only undulations where the even-order Green’s function $\bm{G}^{\text{Dyson}(8)}$ has poles. The fact that the perturbation approximation displays such an abrupt, qualitative change from one order to the next already signals a fundamental failure. Furthermore, the poles of $\bm{G}^{\text{Dyson}(8)}$ tend to occur at HF orbital energy differences (see Fig. 12), and are completely different from the exact poles (superposed in Fig. 14). This is in contrast with other molecular applications of perturbation theory where the eighth order is usually an overkill, yielding near-exact results. For all these reasons, it is concluded that MBGF($n$) is fundamentally not convergent for many low- or high-lying roots.
@@
Figure 15: The third eigenvalue of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and of $\bm{\epsilon}+\bm{\Sigma}^{(n)}(\omega)$ ($2\leq n\leq 9$) as functions of $\omega$ for the BH molecule. The corresponding exact IP obtained from FCI is superposed.
@@
This conclusion does not undermine the utility or rapid convergence of MBGF($n$) for roots falling well within the central bracket bounded by the highest 2h1p and lowest 2p1h energies as well as in other $\omega$ domains (see Sec. V). This bracket typically encloses most principal IPs and EAs (but not core IPs) as well as low-energy satellite poles. Figure 15 is a close-up of the self-energy-versus-$\omega$ plot for the second- through ninth-order MBGF as well as the exact MBGF method. It shows that the intersections of the self-energy and the $\omega$ line systematically approach the exact IP for the HOMO.
IV Infinite partial summations of propagator diagrams
IV.1 Formalisms

### v8

@@ III.1 Formalisms
In most applications, both the Green’s function and self-energy are expanded in perturbation series. In this article, perturbation corrections are denoted by symbols prefixed with $\delta$ with its order given as the parenthesized superscript.
@@
In this article, MBGF($n$) refers to the $n$th-order Feynman–Dyson perturbation approximation to one-particle many-body Green’s function theory, seeking the poles and residues of $\bm{G}^{(n)}$ by solving the inverse Dyson equation with $\bm{\Sigma}^{(n)}$. On the other hand, the exact MBGF is synonymous to FCI as demonstrated in Sec. II. MBGF($n$) is diagrammatically linked Hirata et al. (2017) and thus size-consistent Hirata (2011). Its ab initio computer implementations based on Gaussian-type orbitals Szabo and Ostlund (1982) can therefore be applied to both electrons Hirata et al. (2023) and phonons Qin and Hirata (2020) in solids.
@@
where $\bm{G}(\omega;\lambda)$ and $\bm{\Sigma}(\omega;\lambda)$ are the exact (i.e., FCI) values of the respective quantities for a perturbation-scaled Hamiltonian $\hat{H}=\hat{H}^{(0)}+\lambda\hat{V}^{(1)}$. The zeroth-order Hamiltonian $\hat{H}^{(0)}$ corresponds to the $\bm{G}^{(0)}$ of Eq. (5). Here, we adopt the HF theory as the zeroth order, which implies $\delta{\Sigma}_{pq}^{(1)}={0}$ Hirata et al. (2017).
These $\lambda$-derivatives can be taken either numerically or analytically. From the former, we obtain benchmark data of the perturbation corrections at several low orders Hirata et al. (2017). From the latter, we derive recursions of $\delta\bm{G}^{(n)}$ and $\delta\bm{\Sigma}^{(n)}$ in the style of Rayleigh–Schrödinger perturbation theory, which can then be implemented into a general-order algorithm Hirata et al. (2017). This strategy was used in this study. The recursions also justify the diagrammatic rules through the linked-diagram and irreducible-diagram theorems in a time-independent picture Hirata et al. (2017).
While the time-independent picture is more mathematically transparent and systematically extensible to arbitrarily high orders Hirata et al. (2017), the time-dependent one may be more appealing to our intuition and expedient Dyson (1993); Mattuck (1992). In the latter, for example, the second-order self-energy is stipulated diagrammatically as in Fig. 5. It graphically describes the process in which (i) a mean-field particle (hole) scatters another particle out of its mean-field state, thereby creating a hole, at one time; (ii) all three particles and holes propagate in their respective mean-field potentials, i.e., driven by the mean-field propagator $\bm{G}^{(0)}$; and (iii) the particle-hole pair recombines at another time. The numerical value of $\delta\bm{\Sigma}^{(2)}$, which is related to the probability of this overall process, is the product of the probabilities of the constituent scattering and propagation events summed over all possible times and positions of their occurrences. Consulting with Table 4.3 of Mattuck Mattuck (1992), we can then evaluate it as
@@
where an occupied spinorbital index and the imaginary unit both denoted by “$i$” should be distinguished, and $\langle pq||rs\rangle$ is an antisymmetrized two-electron integral Szabo and Ostlund (1982); Shavitt and Bartlett (2009). Second-order many-body Green’s function method [MBGF(2)] solves the inverse Dyson equation [Eq. (9)] with this $\bm{\Sigma}^{(2)}$. Since the roots of this equation occur at the intersection of $\bm{\epsilon}+\bm{\Sigma}^{(2)}(\omega)$ and $\omega$, they can never be divergent even though the method is perturbative.
@@
in analogy to Eq. (7) and in accordance with the Dyson equations [Eqs. (4) and (4)]. Therefore, as shown in Fig. 7, $\bm{G}^{\text{Dyson}(2)}$ is a bold-line Green’s function (just like the one appearing in Fig. 1), which includes an infinite-order correction through repeated actions of $\bm{\Sigma}^{(2)}$. In this sense, MBGF($n$) is an infinite-order theory for IPs and EAs even for a finite perturbation order $n$.
@@
corresponding to the diagrammatic equation in Fig. 8. It may be noticed that the functional form of $\delta\bm{\Sigma}^{(3)}$ with respect to $\omega$ is different from that of $\delta\bm{\Sigma}^{(2)}$ [Eq. (25)]; poles in $\delta\bm{\Sigma}^{(3)}$ are second order, while $\delta\bm{\Sigma}^{(2)}$ and the exact Green’s function have only first-order poles. This difference has a grave consequence on the roots of the inverse Dyson equation, which we now discuss.
@@ III.2 Numerical results
Figure 9: The third diagonal elements of $\bm{\epsilon}+\bm{\Sigma}^{(2)}(\omega)$ and $\bm{\epsilon}+\bm{\Sigma}^{(3)}(\omega)$ as functions of $\omega$ for the BH molecule.
@@
Figure 10: The third diagonal elements of $\bm{G}^{\text{Dyson}(2)}(\omega)$ and $\bm{G}^{\text{Dyson}(3)}(\omega)$ as functions of $\omega$ for the BH molecule.
@@
In Fig. 9 are plotted diagonal elements of the second- and third-order self-energies, $\epsilon_{3}+\Sigma_{33}^{(2)}$ and $\epsilon_{3}+\Sigma_{33}^{(3)}$, for the third orbital (HOMO) of the BH molecule Hirata et al. (2017). Intersections of these functions with the diagonal $\omega$ line (also drawn) are the roots of the corresponding inverse Dyson equations in the diagonal approximation. See Eq. (13).
@@
The second-order self-energy has qualitatively the same functional form as the exact self-energy (Fig. 3) in that they are both separated by singularities into consecutive $\omega$ brackets, within each of which they are monotonically decreasing and shoulder-like. The singularities of $\bm{\Sigma}^{(2)}$, i.e., the boundaries of the brackets, occur at the two-particle-one-hole (2p1h) ($\epsilon_{a}+\epsilon_{b}-\epsilon_{i}$) and two-hole-one-particle (2h1p) ($\epsilon_{i}+\epsilon_{j}-\epsilon_{a}$) HF orbital energy differences according to Eq. (25). In each bracket, a diagonal element of the second-order self-energy intersects the diagonal $\omega$ line exactly once, just as the exact self-energy does Walter et al. (1984). The second-order inverse Dyson equation has the principal root for the HOMO at around $-0.25\,E_{\text{h}}$ in the central bracket enclosing $\omega=0$, where the self-energy is relatively flat and whose residue is close to unity [see Eq. (14)]. The self-energy also intersects the $\omega$ line in each of the other brackets, typically on a near-vertical part of its shoulder-like curve. This means that these satellite roots coincide with the 2p1h or 2h1p HF orbital energy differences, accounting for little to no electron-correlation effects, and have near zero residues. Nevertheless, MBGF(2) is overall well-behaved.
@@
The third-order self-energy, in contrast, has a qualitatively wrong functional form. It is still separated by the same 2p1h and 2h1p singularities into the identical set of brackets [see Eq. (27)], but within each bracket, $\bm{\Sigma}^{(3)}$ is either concave or convex except in the central bracket, where it is correctly monotonically decreasing. Consequently, in the domain of $\omega$ of Fig. 9, the third-order inverse Dyson equation has only one real root, i.e., the principal root for the HOMO at around $-0.25\,E_{\text{h}}$. The different functional forms of the self-energy between the second and third orders can be easily rationalized by their algebraic definitions, Eqs. (25) and (27). The second-order self-energy (as well as the exact self-energy) has only first-order poles, whereas the third-order self-energy features up to second-order poles.
@@
Figure 11: All six diagonal elements of $\bm{\epsilon}+\bm{\Sigma}^{(3)}(\omega)$ as functions of $\omega$ for the BH molecule. The fourth and fifth elements are degenerate, appearing as one curve. The 2p1h and 2h1p HF orbital energy differences are superposed.
Where have the missing satellite roots of MBGF(3) gone? We believe that they have gone to the complex space, although no attempt has been made to determine them numerically.
@@
From Fig. 11, it can furthermore be observed that real satellite roots of MBGF(3) tend to appear as nearly degenerate pairs or doublets (which is distinct from the exact degeneracy of roots mentioned in the foregoing paragraphs). One of a doublet corresponds to an intersection of a sharply falling part of the self-energy with the $\omega$ line and thus has a positive infinitesimal residue, while the other intersection occurs at an almost vertically rising part of the self-energy, whose residue is negative and thus nonphysical. These intersections almost coincide with bare 2p1h or 2h1p HF orbital energy differences, thus accounting for little to no correlation effects.
Figure 12: The third diagonal elements of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and $\bm{\epsilon}+\bm{\Sigma}^{(n)}(\omega)$ ($n=8,9$) as functions of $\omega$ for the BH molecule. The exact IPs and EAs obtained by FCI are superposed.
@@
In Fig. 12, the odd-order $\bm{\Sigma}^{(9)}$ has a concave or convex shape in each bracket separated by its singularities and, as a result, intersects the $\omega$ line only once in the shown $\omega$ domain. In the same figure, the exact self-energy and exact IPs and EAs are overlaid, indicating that there should be far more intersections than just one in this $\omega$ domain. The missing satellite roots of MBGF(9) are believed to be complex.
The eighth-order self-energy is less problematic in this regard, but it comes with new issues. Its brackets are separated by the 2p1h, 2h1p, 3p2h, 3h2p, 4p3h, 4h3p, 5p4h, and 5h4p HF orbital energy differences. The even-order $\bm{\Sigma}^{(8)}$ is technically monotonically decreasing within each bracket, but it decreases so sharply in many brackets that it consists of near vertical lines, where it intersects the $\omega$ line, as seen in Figs. 12 and 13. Therefore, although MBGF(8) has real roots, most of them are phantom poles with zero residues, occurring somewhere near the midpoint of each bracket.
Consequently, the distribution of the roots of the eighth-order inverse Dyson equation (the intersections of $\bm{\epsilon}+\bm{\Sigma}^{(8)}$ with the $\omega$ line in Fig. 13) matches more closely the mere HF orbital energy differences (the green diamond plots in the same figure) than the exact roots (the red circles). Even though these MBGF(8) roots are not equal to the HF orbital energy differences, they are confined within brackets demarcated by these differences, which form dense manifolds, and tend to agree with them. These satellite roots therefore include little to no electron-correlation effects, not to mention that most of them hardly exist physically because their residues are zero.
These are disheartening results considering that the eighth- or ninth-order perturbation approximations are usually numerically exact in most other physics cases Laidig et al. (1985); Hirata and Bartlett (2000b).
@@
Our numerical tests for other molecules (not shown) have indicated that an odd-order self-energy can generally be concave, convex, or even monotonically increasing within a bracket. Therefore, an odd-order inverse Dyson equation can have real roots (sometimes more than one in a bracket), but the corresponding residues may fall outside the valid range of zero to one.
@@
Figure 14: The third diagonal elements of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and $\bm{\epsilon}+\bm{\Sigma}^{(n)}(\omega)$ ($n=2,4,6,\dots,16$) as functions of $\omega$ for the BH molecule.
One might still wonder if selectively raising only the even perturbation order may avoid these pathologies. Figure 14 shows the evolution of $\bm{\Sigma}^{(2n)}$ as a function of the even perturbation order $2n$. With increasing order, the self-energy becomes more and more vertical and less shoulder-like with the intersections with the $\omega$ line moving away from the correct locations. The corresponding residues will approach zero. Ironically, it appears as though the perturbative self-energies converged at the exact one as the perturbation order is lowered! This figure is the most compelling piece of evidence demonstrating the nonconvergence of the Feynman–Dyson perturbation expansion of the self-energy.
Figure 15: The third eigenvalues of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and $\bm{\epsilon}+\bm{\Sigma}^{(n)}(\omega)$ ($n=2,3,4,\dots,11$) as functions of $\omega$ for the BH molecule. The corresponding exact IP obtained from FCI is superposed.
@@
This conclusion does not undermine the utility or rapid convergence of MBGF($n$) for roots falling well within the central bracket bounded by the highest 2h1p and lowest 2p1h energies as well as in some other $\omega$ domains (see Sec. V.1). This bracket typically encloses most principal IPs and EAs (but not core IPs) as well as low-energy satellite poles. Figure 15 is a close-up of the self-energy-versus-$\omega$ plot for MBGF($n$) ($2\leq n\leq 11$) as well as the exact MBGF. It shows that the intersections of the self-energy and $\omega$ line systematically approach the exact IP for the HOMO.
IV Infinite partial resummations of propagator diagrams
IV.1 Vertex renormalization
To avoid malaise of truncated perturbation approximations, infinite partial resummations of diagrams have often been invoked. One way of doing this is by renormalizing vertexes.
Here, a “vertex” refers to an instantaneous Coulomb interaction denoted by a horizontal (dashed) line in a Hugenholtz–Shavitt–Bartlett diagram and is equal to an anti-symmetrized two-electron integral containing both the direct and exchange contributions Szabo and Ostlund (1982); Shavitt and Bartlett (2009). It should be distinguished from a spacetime point of contact between propagator lines and an interaction line in a Goldstone diagram, each standing for only the direct or exchange contribution, often used in the $GW$ approximations Hedin (1965); Hybertsen and Louie (1986); Aryasetiawan and Gunnarsson (1998); Onida et al. (2002); Vlček (2019); Mejuto-Zaera and Vlček (2022).